QUESTION 10 Big Trail Running Company has started to produce running apparel in addition to the trail running shoes that they have manufactured for years. They feel that a departmental overhead rate would best reflect their overall manufacturing overhead usage. Based on research the following information was gathered for the upcoming year: ​ Machining Department Finishing Department Estimated Manufacturing Overhead by Department $500,000 $400,000 Trail Running Shoes 160,000 machine hours 16,000 direct labor hours Running Apparel 40,000 machine hours 64,000 direct labor hours Manufacturing overhead is driven by machine hours for the machining department and direct labor hours for the finishing department. At the end of the year, the following information was gathered related to the production of the trail running shoes and running apparel: ​ Machining Department Finishing Department Trail Running Shoes 162,000 hours 15,500 hours Running Apparel 37,000 hours 65,000 hours How much manufacturing overhead will be allocated to the trail running shoes? (Round any intermediary calculations to the nearest cent and your final answer to the nearest dollar.) $482,500 $900,000 $484,054 $225,000
Answer #1
Departmental overhead rate:
Machining Department 2.50 =500000/(160000+40000)
Finishing Department 5.00 =400000/(16000+64000)
Trail Running Shoes
Manufacturing overhead allocated:
Machining Department 405000 =162000*2.50
Finishing Department 77500 =15500*5.00
Manufacturing overhead allocated 482500
Option A $482,500 is correct